# Geodesic Distance Between Graphs: A Spectral Metric for Assessing the Stability of Graph Neural Networks

### v2

@@ Abstract
This paper introduces a spectral framework for assessing the generalization of Graph Neural Networks (GNNs) through a new Graph Geodesic Distance (GGD) metric. For two different graphs with the same number of nodes, our framework leverages a spectral graph matching procedure to find node correspondence so that the geodesic distance between them can be subsequently computed by solving a generalized eigenvalue problem associated with their Laplacian matrices. For graphs of different sizes, a resistance-based spectral graph coarsening scheme is introduced to reduce the size of the larger graph while preserving the original spectral properties. We show that the proposed GGD metric can effectively quantify dissimilarities between two graphs by encapsulating their differences in key structural (spectral) properties, such as effective resistances between nodes, cuts, the mixing time of random walks, etc. Through extensive experiments comparing with the state-of-the-art metrics, such as the latest Tree-Mover’s Distance (TMD) metric, the proposed GGD metric shows significantly improved performance for graph classification and stability evaluation of GNNs, especially when only partial node features are available.
@@ 1 Introduction
In the era of big data, comparison and distinction between data points are important tasks. A graph is a specific type of data structure that represents the connections between a group of nodes or agents. Comparing two graphs often involves using a pairwise distance measure, where a small distance indicates a high structural similarity and vice versa. To understand the generalization between distribution shifts, it is important to use an appropriate measure of divergence between data distributions, both theoretically and experimentally (Chuang et al., 2020). Determining suitable distance metrics for non-Euclidean data entities, like graphs with or without node attributes, which are fundamental to many graph learning methods such as graph neural networks (GNNs), remains a significant challenge, even though distance metrics for data points in Euclidean space are readily available. The need to develop new analytical techniques that allow the visualization, comparison, and understanding of different graphs has led to a rich field of research study (Haslbeck & Waldorp, 2018). This study dives into the exploration of a novel framework for computing geometric distances between graphs, which can be immediately leveraged for many graph-based machine learning (ML) tasks, such as graph classification or the stability evaluation of GNNs.
@@
Many distance metrics for comparing graphs have previously been proposed (Borgwardt et al., 2020). Some of them are merely based on graph local structures (Tam & Dunson, 2022; Haussler et al., 1999; Xu et al., 2013; Zhu et al., 2020; Fernández & Valiente, 2001; Bunke & Shearer, 1998), whereas others exploit both graph structural properties and node attributes (Shervashidze et al., 2011; Morris et al., 2019). For example, the Graph Edit Distance (GED) has been proposed to measure the distance between graphs considering the number of changes needed to match one graph to another (Sanfeliu & Fu, 1983; Gao et al., 2010; Li et al., 2017); Distance metrics based on the graph kernel have also been investigated (Shervashidze et al., 2011; Vishwanathan et al., 2010), such as the Wasserstein Weisfeiler-Leman metric (WWL) (Morris et al., 2019) and the Gromov–Wasserstein metric (Mémoli, 2011), which allow computing graph distances based on low-dimensional graph representations or optimal transport (OT) (Titouan et al., 2019; Chapel et al., 2020), leading to the development of the state-of-the-art graph distance metric called TMD (Chuang & Jegelka, 2022).
@@
However, the existing graph distance metrics have distinct limitations. For example, the GED metric can capture local node or edge changes but struggles with global perturbations (Sanfeliu & Fu, 1983; Gao et al., 2010; Li et al., 2017); the WWL and TMD metrics heavily rely on node features (attributes) for calculating the distance between graphs, leading to degraded performance when only partial node features are available (Rossi et al., 2022; Chen et al., 2022).
To address the limitations of prior methods, we propose a Graph Geodesic Distance (GGD) metric that can be computed through the following steps: for two graphs of the same size (with $n$ nodes), (1) a spectral graph matching procedure is exploited for finding the node correspondence between two graphs, and (2) a geodesic distance metric is defined on the cone of $n\times n$ modified graph Laplacians that can be regarded as a Riemannian manifold (Lim et al., 2019). For graphs of different sizes, we introduce a resistance-based spectral graph coarsening scheme to transform the larger graph into a smaller one suitable for the GGD evaluation. We show that the proposed GGD metric can theoretically capture key structural (spectral) dissimilarities between two graphs, such as mismatches in Laplacian eigenvalues/eigenvectors, cuts, effective-resistance distances, etc.
@@
One distinct advantage of the proposed GGD metric is its capability to compute distances between graphs based on their spectral (structural) properties, while including node feature information into our framework can further improve the accuracy. Therefore, GGD is suitable for analyzing many real-world graphs that may only have partial or even no node features. Moreover, the proposed framework for computing GGDs is more computationally efficient than existing OT-based metrics, such as the TMD metric.
@@
Our empirical results show that GGD can effectively measure the dissimilarities between graphs: (1) support vector classifiers (SVC) using GGDs perform competitively with state-of-the-art GNN models and graph kernels on graph classification benchmarks; (2) we demonstrate that the GGD metric allows us to quantify the stability of GNN models for graph classification tasks by checking whether two graphs with a small GGD will lead to a significant dissimilarity in the GNN output embeddings. We also show that the GGD metric has a better correlation with established GNN outputs compared to the state-of-the-art TMD metric (Chuang & Jegelka, 2022) when only partial node features are available: up to a $10\%$ accuracy gain and a 9×9$\times$9 × runtime speedup have been achieved in various graph classification tasks.
@@ 2 Existing Graph Distance Metrics
Graph Edit Distance (GED) For non-attributed graph data, a common and simple distance metric is GED. (Sanfeliu & Fu, 1983; Gao et al., 2010). Given a set of graph edit operations, also known as elementary graph operations, the GED between two graphs $G_{1}$ and $G_{2}$, written as $\text{GED}(G_{1},G_{2})$, can be defined as:
@@
Tree Mover’s Distance (TMD) TMD is a pseudometric for measuring distances between simple graphs, extending the concept of WWL to multisets of tree structures (Chuang & Jegelka, 2022). By progressively adding neighboring nodes to the previous node at each level, we obtain the computation tree of a node. These tree structures are crucial in graph analysis (Weisfeiler & Leman, 1968; Pearson, 1905) and graph kernels (Ramon & Gärtner, 2003; Shervashidze et al., 2011). TMD uses hierarchical optimal transport (HOT) to analyze these computational trees from input graphs. For a graph $G=(V,E)$ with node features $f_{v}\in\mathbb{R}^{s}$ for node $v\in V$, let $T_{v}^{1}=v$, and $T_{v}^{L}$ be the depth-L computation tree of node $v$. The multiset of these trees for $G$ is $T_{G}^{L}=\{T_{v}^{L}\}_{v\in V}$. The number and shape of trees must match to calculate optimal transport between two multisets of trees. If multisets are uneven, they are augmented with blank nodes. For multisets $T_{p}$ and $T_{q}$, the augmenting function $\sigma$ adds blank trees to equalize their sizes. A blank tree $T_{\mathbb{O}}$ has a single node with a zero vector feature $\mathbb{O}_{p}\in\mathbb{R}^{s}$:
$\sigma:(T_{p},T_{q})\rightarrow\left(T_{p}\cup T_{\mathbb{O}}^{\max(|T_{q}|-|T_{p}|,0)},T_{q}\cup T_{\mathbb{O}}^{\max(|T_{p}|-|T_{q}|,0)}\right).$ (2)
Let $X=\left\{x_{i}\right\}_{i=1}^{k}$ and $Y=\left\{y_{i}\right\}_{j=1}^{k}$ be two data multisets and $C\in\mathbb{R}^{k\times k}$ be the transportation cost for each data pair: $C_{ij}=d\left(x_{i},y_{j}\right)$, where $d$ is the distance between $x_{i}$ and $y_{j}$. The unnormalized Optimal Transport between $X$ and $Y$ can be defined as:
$\mathrm{OT}_{d}(X,Y):=\min_{\gamma\in\Gamma(X,Y)}\langle C,\gamma\rangle,\quad\Gamma(X,Y)=\left\{\gamma\in\mathbb{R}_{+}^{m\times m}\mid\gamma\mathbb{1}_{m}=\gamma^{\top}\mathbb{1}_{m}=\mathbb{1}_{m}\right\}.$ (3)
Here $\Gamma$ is the set of transportation plans that satisfies the flow constrain $\gamma\mathbb{1}_{m}=\gamma^{\top}\mathbb{1}_{m}=\mathbb{1}_{m}$. (Chuang & Jegelka, 2022).
The distance between two trees $T_{p}$ and $T_{q}$ with roots $r_{p}$ and $r_{q}$ is defined recursively:
$\text{TD}_{w}\left(T_{p},T_{q}\right):=\begin{cases}\left\|f_{r_{p}}-f_{r_{q}}\right\|+w(L)\cdot\mathrm{OT}_{\mathrm{TD}_{w}}\left(\sigma\left(T_{r_{p}},T_{r_{q}}\right)\right),&\text{if }L>1\\ \left\|f_{r_{p}}-f_{r_{q}}\right\|,&\text{otherwise}\end{cases}$ (4)
@@
Subsequently, the concept of distance from individual trees is enlarged to entire graphs. For graphs $G_{1}$ and $G_{2}$, with multisets $\mathbf{T}_{G_{1}}^{L}$ and $\mathbf{T}_{G_{2}}^{L}$ of depth-L computation trees, the Tree Mover’s Distance is:
@@
3 GGD: A Geodesic Distance Metric for Graphs
@@
To overcome the limitations of previous OT-based graph distance metrics (Chuang & Jegelka, 2022), we aim to develop a new metric that effectively captures topological differences between graphs. One way to represent a simple graph is through its Laplacian matrix, which is a Symmetric Positive Semidefinite (SPSD) matrix. Graph representation using adjacency and Laplacian matrices is briefly discussed in Appendix A.1. Adding a small positive value to each diagonal element allows us to transform the original Laplacian matrix into a Symmetric Positive Definite (SPD) matrix, which is referred to as the Modified Laplacian Matrix in this work. In Appendix A.6, we describe the effect of this small value on the GGD calculation. We can then consider the cone of such modified Laplacian matrices as a natural Riemannian manifold (Lim et al., 2019), where each modified Laplacian, having the same dimensions (same number of rows/columns), can be regarded as a data point on this Riemannian manifold (Vemulapalli & Jacobs, 2015; Pennec et al., 2006). Details about the Riemannian manifold are provided in Appendix A.2. Finally, the geodesic distance between two graphs can be defined as the shortest path distance on the Riemannian manifold, assuming their node correspondence is known in advance. This approach is more appropriate than directly comparing the graphs in Euclidean space (Lim et al., 2019; Crane et al., 2020; Huang et al., 2015). We will later demonstrate (Section 4.3) that such a geodesic distance metric can effectively capture structural (spectral) mismatches between graphs.
A two-phase spectral framework for computing GGDs
@@
Before computing GGDs, it is necessary to establish the node-to-node correspondence between two graphs. This can be achieved by leveraging existing graph-matching techniques (Livi & Rizzi, 2013; Emmert-Streib et al., 2016; Caetano et al., 2009). In this work, we will leverage a recent spectral graph matching framework that has been shown to recover accurate matching with high probability (Fan et al., 2020). The proposed GGD metric can be computed in the following two phases. Phase 1 consists of a spectral graph matching step, using combinatorial optimization with the eigenvalues/eigenvectors of the graph adjacency matrices to identify the approximate node-to-node correspondence. Phase 2 computes the GGD between the modified Laplacian matrices of the matched graphs by exploiting generalized eigenvalues. The proposed GGD metric differs from previous OT-based graph distance metrics in its ability to accurately represent structural discrepancies between graphs, enabling us to uncover the topological variations between them more effectively. Since only the graph Laplacian (adjacency) matrix is required to calculate the GGD, our metric can even work effectively for graphs without node feature information.
A motivating example
Let’s consider a simple graph $G_{1}$, characterized by an almost ring-like topology, as shown in Figure 1. We also create two other graphs $G_{2}$ and $G_{3}$ by inserting an extra edge into $G_{1}$ in different ways. Note that the additional edge in $G_{3}$ will have a greater impact on $G_{1}$’s global structure since it connects two further nodes.
@@
Table 1: Distance between graphs with simple perturbations.
{tblr}
roweven = c, row3 = c, row5 = c, cell11 = r=2, cell12 = c=4c, hline1,3,6 = -, hline2 = 2-5, Graph pairs & Distance metrics (Normalized)
GGD TMD with NF, L = 4 TMD without NF, L = 4 GED
$G_{1},G_{2}$ 0.623 0.689 0.970 1
$G_{1},G_{3}$ 0.855 0.711 1 1
$G_{2},G_{3}$ 1 1 0.333 1
@@
We compute the normalized distances (the largest distance always equals one) between the aforementioned three graphs using different metrics (GED, TMD, and GGD) and report the results in Table 1. As observed, $G_{2}$ and $G_{3}$ have distances similar to $G_{1}$ when the TMD metric is adopted without using node features (NFs). On the other hand, the TMD metric can produce similar results as the proposed GGD metric when node features are fully utilized. Not surprisingly, the GED always produces the same distances since only one edge has been added. The above results imply that the GED and TMD (without using NFs) metrics may not properly capture the dissimilarities in the structural (spectral) properties of the graphs.

### v3

# A Spectral Framework for Evaluating Geodesic Distances Between Graphs

@@ Abstract
This paper presents a spectral framework for quantifying the differentiation between graph data samples by introducing a novel metric named Graph Geodesic Distance (GGD). For two different graphs with the same number of nodes, our framework leverages a spectral graph matching procedure to find node correspondence so that the geodesic distance between them can be subsequently computed by solving a generalized eigenvalue problem associated with their Laplacian matrices. For graphs of different sizes, a resistance-based spectral graph coarsening scheme is introduced to reduce the size of the larger graph while preserving the original spectral properties. We show that the proposed GGD metric can effectively quantify dissimilarities between two graphs by encapsulating their differences in key structural (spectral) properties, such as effective resistances between nodes, cuts, and the mixing time of random walks. Through extensive experiments comparing with state-of-the-art metrics, such as the latest Tree-Mover’s Distance (TMD), the proposed GGD metric demonstrates significantly improved performance for graph classification, particularly when only partial node features are available. Furthermore, we extend the application of GGD beyond graph classification to stability analysis of GNNs and the quantification of distances between datasets, highlighting its versatility in broader machine learning contexts.
@@ 1 Introduction
In the era of big data, comparison and distinction between data points are important tasks. A graph is a specific type of data structure that represents the connections between a group of nodes. Comparing two graphs often involves using a pairwise distance measure, where a small distance indicates a high structural similarity and vice versa. To understand the generalization between distribution shifts, it is important to use an appropriate measure of divergence between data distributions, both theoretically and experimentally (chuang2020estimating, ). Determining suitable distance metrics for non-Euclidean data, such as graphs with or without node attributes, remains challenging. These metrics are fundamental to many graph learning methods, such as graph neural networks (GNNs), but are not as readily available as those for Euclidean space. The need to develop new analytical techniques that allow the visualization, comparison, and understanding of different graphs has led to a rich field of research study (haslbeck2018well, ). This study dives into the exploration of a novel framework for computing geometric distances between graphs, which can be immediately leveraged for many graph-based machine learning (ML) tasks, such as graph classification, dataset distance, or the stability evaluation of GNNs.
@@
Many distance metrics for comparing graphs have previously been proposed (MAL, ). Some of them are based only on graph local structures (tam2022multiscale, ; haussler1999convolution, ; xu2013distance, ; zhu2020fewer, ; fernandez2001graph, ; bunke1998graph, ), whereas others exploit both graph structural properties and node attributes (shervashidze2011weisfeiler, ; morris2019weisfeiler, ). For example, the Graph Edit Distance (GED) has been proposed to measure the distance between graphs considering the number of changes needed to match one graph to another (sanfeliu1983distancenew, ; gao2010survey, ; li2017comparative, ); Distance metrics based on the graph kernel have also been investigated (shervashidze2011weisfeiler, ; vishwanathan2010graph, ), such as the Wasserstein Weisfeiler-Leman metric (WWL) (morris2019weisfeiler, ) and the Gromov–Wasserstein metric (memoli2011gromov, ), which allow computing graph distances based on low-dimensional graph representations or optimal transport (OT) (titouan2019optimal, ; chapel2020partial, ), leading to the development of the state-of-the-art graph distance metric called TMD (chuang2022tree, ).
@@
However, existing graph distance metrics have notable limitations. For example, the GED metric can capture local node or edge changes but struggles with global perturbations (sanfeliu1983distancenew, ; gao2010survey, ; li2017comparative, ); the WWL and TMD metrics heavily rely on node features (attributes) for calculating the distance between graphs, leading to degraded performance when only partial node features are available (rossi2022unreasonable, ; chen2020learning, ).
To address these limitations, we propose the Graph Geodesic Distance (GGD) metric—an end-to-end graph comparison framework that leverages spectral graph theory, structure-preserving coarsening, and Riemannian geometry to compute meaningful distances between graphs.
To address these limitations of prior methods, we propose the Graph Geodesic Distance (GGD) metric, a novel framework that combines spectral graph theory and Riemannian geometry to quantify topological distances between graphs. Unlike prior SPD-based works (lim2019geometric, ), our method operates directly on graph inputs, using spectral graph matching to establish node correspondence before embedding graphs into a Riemannian manifold of modified Laplacian matrices. This allows GGD to effectively capture key structural (spectral) dissimilarities between two graphs, such as mismatches in Laplacian eigenvalues/eigenvectors, cuts, effective-resistance distances, etc.
@@
One distinct advantage of the proposed GGD metric is its capability to compute distances between graphs based on their spectral (structural) properties, while including node features can further improve its accuracy. This makes GGD suitable for analyzing real-world graphs with partial or no node features. Moreover, the proposed framework for computing GGDs is more computationally efficient than existing OT-based metrics, such as the TMD metric.
@@
Our empirical results show that GGD can effectively measure the dissimilarities between graphs: (1) support vector classifiers (SVC) using GGDs perform competitively with state-of-the-art GNN models and graph kernels on graph classification benchmarks; (2) we demonstrate that the GGD metric allows us to quantify the stability of GNN models for graph classification tasks by checking whether two graphs with a small GGD will lead to a significant dissimilarity in the GNN output embeddings; (3) quantify distance between datasets to evaluate the transferability of domain knowledge. We also show that the GGD metric has a better correlation with established GNN outputs compared to the state-of-the-art TMD metric (chuang2022tree, ) when only partial node features are available: up to a $10\%$ accuracy gain and a $9\times$ runtime speedup have been achieved in various graph classification tasks.
@@ 2 Existing Graph Distance Metrics
Graph Edit Distance (GED) For non-attributed graph data, a common and simple distance metric is GED. (sanfeliu1983distancenew, ; gao2010survey, ). Given a set of graph edit operations, also known as elementary graph operations, the GED between two graphs $G_{1}$ and $G_{2}$, written as $\text{GED}(G_{1},G_{2})$, can be defined as:
@@
Tree Mover’s Distance (TMD) TMD is a pseudometric for measuring distances between simple graphs, extending the concept of WWL to multisets of tree structures (chuang2022tree, ). By progressively adding neighboring nodes to the previous node at each level, we obtain the computation tree of a node. These tree structures are crucial in graph analysis (weisfeiler1968reduction, ; pearson1905problem, ) and graph kernels (ramon2003expressivity, ; shervashidze2011weisfeiler, ). TMD uses hierarchical optimal transport (HOT) to analyze these computational trees from input graphs. For a graph $G=(V,E)$ with node features $f_{v}\in\mathbb{R}^{s}$ for node $v\in V$, let $T_{v}^{1}=v$, and $T_{v}^{L}$ be the depth-L computation tree of node $v$. The multiset of these trees for $G$ is $T_{G}^{L}=\{T_{v}^{L}\}_{v\in V}$. The number and shape of trees must match to calculate optimal transport between two multisets of trees. If multisets are uneven, they are augmented with blank nodes. For multisets $T_{p}$ and $T_{q}$, the augmenting function $\sigma$ adds blank trees to equalize their sizes. A blank tree $T_{\mathbb{O}}$ has a single node with a zero vector feature $\mathbb{O}_{p}\in\mathbb{R}^{s}$:
$\sigma(T_{p},T_{q})\rightarrow\Big{(}T_{p}\cup T_{\mathbb{O}}^{\max(|T_{q}|-|T_{p}|,0)},\;T_{q}\cup T_{\mathbb{O}}^{\max(|T_{p}|-|T_{q}|,0)}\Big{)}$ (2)
Let $X=\left\{x_{i}\right\}_{i=1}^{k}$ and $Y=\left\{y_{i}\right\}_{j=1}^{k}$ be two data multisets and $C\in\mathbb{R}^{k\times k}$ be the transportation cost for each data pair: $C_{ij}=d\left(x_{i},y_{j}\right)$, where $d$ is the distance between $x_{i}$ and $y_{j}$. The unnormalized Optimal Transport between $X$ and $Y$ is defined as follows:
$\begin{array}[]{l}\mathrm{OT}_{d}(X,Y):=\min_{\gamma\in\Gamma(X,Y)}\langle C,\gamma\rangle\\ \Gamma(X,Y)=\left\{\gamma\in\mathbb{R}_{+}^{m\times m}\mid\gamma\mathbbold{1}_{m}=\gamma^{\top}\mathbbold{1}_{m}=\mathbbold{1}_{m}\right\}.\end{array}$ (3)
Here $\Gamma$ is the set of transportation plans that satisfies the flow constrain $\gamma\mathbbold{1}_{m}=\gamma^{\top}\mathbbold{1}_{m}=\mathbbold{1}_{m}$. (chuang2022tree, ). Now, the distance between two trees $T_{p}$ and $T_{q}$ with roots $r_{p}$ and $r_{q}$ is defined recursively:
[…]
@@
Subsequently, the concept of distance from individual trees is extended to entire graphs. For graphs $G_{1}$ and $G_{2}$, with multisets $\mathbf{T}_{G_{1}}^{L}$ and $\mathbf{T}_{G_{2}}^{L}$ of depth-L computation trees, the Tree Mover’s Distance is:
@@
3 GGD: Geodesic Distance Between Graphs
@@
One way to represent a simple connected graph is through its Laplacian matrix, which is a Symmetric Positive Semidefinite (SPSD) matrix. Graph representation using adjacency and Laplacian matrices is briefly discussed in Appendix A.2. Adding a small positive value to each diagonal element allows us to transform the original Laplacian matrix into a Symmetric Positive Definite (SPD) matrix, which is referred to as the Modified Laplacian Matrix in this work. In Appendix A.9, we describe the effect of this small value on the GGD calculation. We can then consider the cone of such modified Laplacian matrices as a natural Riemannian manifold (lim2019geometric, ), where each modified Laplacian, having the same dimensions (same number of rows/columns), can be regarded as a data point on this Riemannian manifold (vemulapalli2015riemannian, ; pennec2006riemannian, ). Details about the Riemannian manifold are provided in Appendix A.3. The geodesic distance is defined as the shortest path on the Riemannian manifold, providing a more appropriate comparison than Euclidean space (lim2019geometric, ; crane2020survey, ; huang2015log, ). We will later demonstrate (Section 4.3) that such a geodesic distance metric can effectively capture structural (spectral) mismatches between graphs.
Figure 1: A high-level illustration of the GGD pipeline, including spectral graph coarsening, (Phase 1) spectral graph matching, and (Phase 2) geodesic distance calculation for the Riemannian manifold of modified Laplacian matrices.
@@
A two-phase spectral framework for computing GGDs Before computing GGDs, it is necessary to establish the node-to-node correspondence between two graphs. This can be achieved by leveraging existing graph-matching techniques (livi2013graph, ; emmert2016fifty, ; caetano2009learning, ). The proposed GGD metric can be computed in the following two phases. Phase 1 consists of a spectral graph matching step, using combinatorial optimization with the eigenvalues/eigenvectors of the graph adjacency matrices to identify the approximate node-to-node correspondence. Phase 2 computes the GGD between the modified Laplacian matrices of the matched graphs by exploiting generalized eigenvalues. A high-level overview of the pipeline is illustrated in Figure 1, and a detailed algorithmic flow is provided in Appendix A.1 to ensure a clear understanding of the process. The proposed GGD metric differs from previous OT-based graph distance metrics in its ability to accurately represent structural discrepancies between graphs, enabling us to uncover the topological variations between them more effectively.
@@
Table 1: Normalized distance between graphs with simple perturbations.
Graph
Pairs GGD
NF, L = 4 TMD w/o
NF, L = 4 GED
$G_{1},G_{2}$ 0.623 0.689 0.970 1.000
$G_{1},G_{3}$ 0.855 0.711 1.000 1.000
$G_{2},G_{3}$ 1.000 1.000 0.333 1.000
A motivating example Let’s consider a simple graph $G_{1}$, characterized by an almost ring-like topology, as shown in Figure 2. We also create two other graphs $G_{2}$ and $G_{3}$ by inserting an extra edge into $G_{1}$ in different ways. Note that the additional edge in $G_{3}$ will have a greater impact on $G_{1}$’s global structure since it connects two further nodes.
@@
We compute the normalized distances (the largest distance always equals one) between the aforementioned three graphs using different metrics (GED, TMD, and GGD) and report the results in Table 3. As observed, $G_{2}$ and $G_{3}$ have distances similar to $G_{1}$ when the TMD metric is adopted without using node features (NFs). On the other hand, the TMD metric can produce similar results as the proposed GGD metric when node features are fully utilized. Not surprisingly, the GED always produces the same distances since only one edge has been added. The above results imply that the GED and TMD (without using NFs) metrics may not properly capture the dissimilarities in the spectral properties of the graphs.